Two swimmers leave point A on one bank of the river to reach point B lying right across on the other bank. One of them crosses the river along the straight line AB while the other swims at right angles to the stream and then walks the distance that he has been carried away by the stream to get to point B. What was the velocity u of his walking if both swimmers reached the destination simultaneously? The stream velocity vo = 2.0 km/hour and the velocity V of each swimmer with respect to water equals 2.5 km per hour.




Grade:12

1 Answers

Swapnil Saxena
12 years ago

The answer is 3km/h.

Assume the width of the river = d

The time taken by the swimmer moving across AB

V(horizontal Com.)= (2.52-22)1/2=(1.5km/h)

t= d/1.5

The time taken by the swimmer swimming at right angles

t= (d/2.5)(time taken to cross the river)+(2 d/(2.5v))(time taken to cross remaining diatnce on foot , v is the veolcity on ground)

Put both of  time equal and solve them to get the answer 3